# Reconfigurable dissipative entanglement between many spin ensembles: from robust quantum sensing to many-body state engineering

### v1

@@ V.4 Tunability and string order for arbitrary spin-$S$
Note, crucially, that for any finite value of $\Delta_{e}/\chi$ and $\Delta_{b}/\chi$, $L_{\rm corr}$ remains finite and $\mathcal{O}^{(z)}_{\rm string}(\pi)$ is always non-zero (see Fig. 8(a)). This result indicates that all the states in Eq. (33) have SPT order and belong to the same phase as the spin-1 AKLT state. This agrees with Ref. 80, which argued that the parameterized MPS in Eq. (33) with $S=1/2$ serves as a path smoothly connecting spin-$1/2$ dimer and spin-$1$ AKLT states. This type of SPT order is due to the emergent SO(3) symmetry in the large-$\Omega$ limit indicated by Eq. (32). For a finite Rabi drive $\Omega$, strictly speaking, there is no true SPT order, as the string order parameter will decay to zero at infinite distance. However, for any finite system size, one can always increase $\Omega$ to realize an approximation of SPT order, where the the string order parameter remains non-zero over a distance much larger than system size (see App. D).
As we have emphasized, a key virtue of our setup is its extreme flexibility. We now show that the above construction with $S=1/2$ directly generalizes to arbitrary spin-$S$ ensembles on each site. One can show that the correlation length $L_{\rm corr}$ remains finite for the general case, i.e., all the states in the large-$\Omega$ limit lie in the same phase. We then perform an analytical calculation of the string order parameter for arbitrary spin $S$, similar to Refs. 77, 78. As shown in App. D, it is possible to separate the dependence of system parameters $\Delta_{e},\Delta_{b},\chi$ and the dependence of the angle $\varphi$,
@@
Notice that we have $\mathcal{O}^{(z)}_{\rm string}(\pi)=0$ for integer $S$ (no SPT order), and $\mathcal{O}^{(z)}_{\rm string}(\pi)>0$ for half-integer $S$ (SPT order). This result agrees with the discussions of spin-$2S$ AKLT states in Ref. 27. We show the distinct dependence of the angle $\varphi$ for different spin $S$ in Fig. 8(b).
@@ VI Conclusion and outlook
The results presented here open several avenues for further exploration. On the metrological side, distributed quantum sensing [81] represents a particularly appealing application. While explicitly analyzed for two- and four-ensemble configurations, our approach can be scaled to larger networks, potentially enabling spatially resolved field detection and improved sensitivity scaling with system size. On the state-engineering side, it is intriguing to explore dissipative stabilization schemes for more complex many-body states. For example, our approach suggests a natural route toward realizing higher-spin AKLT states by further increasing the tunability of the setup by adding additional Hamiltonian terms that have the form of rank-$1$ spherical tensors. In addition, the stabilized states in our scheme bear a clear resemblance to quantum spin liquids [82], both consisting of superpositions of singlets with zero total angular momentum. This analogy suggests that our framework might be extendable to the dissipative preparation of spin-liquid states.
@@ Appendix A Analytic calculation in the total angular momentum basis
This result is directly relevant for the existence of the unitary transformation in Eq. (22). Since $B_{J,m}$ and $C_{J,m}$ are both matrix elements of spherical tensors with $k=1$ and $q=0$, they share the same $m$ dependence due to the Wigner-Eckart theorem. This is the reason why the phases $\theta_{J}$ in Eq. (22) have no $m$ dependence.
@@
Appendix B Lower bound of relaxation time scale
Here we derive a lower bound of the relaxation time scale to a pure steady state $|\psi_{\rm ss}^{(L)}\rangle$ based on Refs. 45, 46. We define relaxation time $t_{\rm ss}$ such that the fidelity with the steady state $F(t)=\mathrm{tr}\Big(\hat{\rho}_{\rm ss}\,\hat{\rho}(t)\Big)$ satisfies $F(t)>1-\epsilon$ for $t>t_{\rm ss}$, where $\epsilon$ is a small positive number, and $\hat{\rho}_{\rm ss}=|\psi_{\rm ss}^{(L)}\rangle\langle\psi_{\rm ss}^{(L)}|$. A lower bound of $t_{\rm ss}$ is given as follows,
@@ Appendix B Lower bound of relaxation time scale
In the case of two spin ensembles ($L=2$), we can further relate the lower bound of $t_{\rm ss}$ to the steady-state spin squeezing, since $\langle\hat{S}_{1}^{z}+\hat{S}_{2}^{z}\rangle_{\rm ss}=\sum_{J=0}^{2S}|c_{J,-J}|^{2}(-J)$, and $\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}=\frac{1}{2}\sum_{J=0}^{2S}|c_{J,-J}|^{2}J$. So we have
@@
As we shown in Fig. 3(b), the optimal $t_{\rm ss}$ obtained numerically is only a constant factor away from this lower bound for two spin ensembles.
@@ D.1 MPS representation of 1D chain
One can also conclude that the matrix $\mathbf{A}^{Jm}$ is a spherical tensor $\hat{T}^{(k)}_{q}$ with $k=J$ and $q=m$ based on the Wigner-Eckart theorem, which is acting on the bond dimension (spin-$S$ particle), i.e., $\sum_{\alpha,\beta}[\mathbf{A}^{Jm}]_{\alpha\beta}|S,\alpha\rangle\langle S,\beta|$. The same structure is used in Ref. 70 to define the valence bond solid (VBS) state.
@@ D.2 String order parameter
Similar to the calculation in Refs. 77, 78, here we would like to evaluate the string order parameter for Eq. (79) in an infinite 1D chain,
@@ D.3 Correlation length
Here, we consider the case of $S=1/2$ and set $\varphi=\pi$. We also set $\Delta_{e}=0$ and $\Delta_{b}=\sqrt{2}\,\chi$ (parameters for spin-$1$ AKLT state in the large-Rabi-drive limit). In Fig. 9(a), a numerical calculation of $L_{\rm string}$ shows that $L_{\rm string}\propto(\Omega/\chi)^{2}$. When $L_{\rm string}\gg L$, we effectively realize the SPT phase since we cannot tell the difference in a finite-size system. One can also interpret $L/L_{\rm string}$ as a measure of the fidelity to the spin-$1$ AKLT state, as shown in Fig. 9(b). As we increase $L$, we find that $L/L_{\rm string}$ and the fidelity collapse into a single line. For example, $90\%$ fidelity corresponds to $L/L_{\rm string}\approx 0.2$. Based on the discussions in the main text, we have the relaxation time scale $t_{\rm ss}\propto(\Omega/\chi)^{2}$, leading to $L_{\rm string}\propto t_{\rm ss}$. This result indicates that the AKLT relaxation time scale $t_{\rm AKLT}\propto L$ if fixing the same fidelity to the spin-$1$ AKLT state.
@@ F.2 Many spin ensembles
Assuming the Lindbladian is diagonalizable, we can extend the uniqueness of the steady-state solution to the case of $\chi\neq\Gamma$ if the permutation symmetry between ensembles is completely broken. We also numerically confirm the uniqueness for the case of $\chi\neq\Gamma$ for small system sizes (up to $L=8$ for $S=1/2$).
@@ G.1 Raman-coupled spin-exchange interactions
One way to implement the chiral spin-exchange couplings is to engineer Raman-coupled spin-exchange interactions similar to the setup in Ref. [22]. As shown in Fig. 12(a), we define $|\uparrow\rangle$ and $|\downarrow\rangle$ states in the ground manifold with transition frequency $\omega_{0}$, and excited state $|e\rangle$ with transition frequency $\omega_{e}$ with respect to $|\uparrow\rangle$ state. We then apply a magnetic-field gradient along the cavity such that the transition frequency between $|\uparrow\rangle$ and $|\downarrow\rangle$ states becomes $\omega_{0,l}=\omega_{0}+(l-1)(\Delta\omega_{0})$. To engineer chiral spin-exchange couplings between nearest-neighbor ensembles (purple arrows), we apply a two-tone drive to transition from $|\uparrow\rangle$ to $|e\rangle$ off-resonantly (solid and dashed arrows). The two tones have Rabi frequency $\Omega^{\prime}$ and $i\Omega^{\prime}$, and drive frequency $\omega_{d}$ and $\omega_{d}-\Delta\omega_{0}$ respectively. We still use the same detuning $\delta_{e}=\omega_{d}-\omega_{e}$ to the excited state $|e\rangle$ assuming $\delta_{e}$ much larger than Zeeman shifts. The cavity mode couples the transition from $|\downarrow\rangle$ to $|e\rangle$ with cavity resonant frequency $\omega_{c}$. After adiabatic elimination of the excited states ($\delta_{e}\gg\Omega^{\prime}$), the Hamiltonian is given by
@@
The second term in Eq. (120) is the chiral spin-exchange coupling between nearest-neighbor ensembles, which is due to frequency selection via the two-tone frequency difference $\Delta_{0}$. Other chiral spin-exchange coupling terms can be engineered by two-tone frequency differences $2\Delta_{0},3\Delta_{0},\cdots$. One can also cancel the first term in Eq. (120) using a single-tone drive that generates $\sum_{l}\hat{S}^{+}_{l}\hat{S}^{-}_{l}$ with opposite $\delta_{c}$.
@@ G.2 Chiral quantum networks
Another way to implement the chiral spin-exchange couplings is to couple each spin ensemble with an optical cavity, and then connect all these cavities by a single-directional waveguide as shown in Fig. 12(b). Based on Refs. [19, 21], the Lindblad master equation of this system can be written as
@@
$\hat{H}_{l}=\Omega\hat{S}^{x}_{l}+\delta_{l}\hat{S}^{z}_{l}+\mathcal{G}(\hat{a}^{{\dagger}}_{l}\hat{S}^{-}_{l}+\hat{S}^{+}_{l}\hat{a}_{l}),$ (122)
@@
where $2\mathcal{G}$ is the single-atom vacuum Rabi splitting of the cavity. When $\kappa\gg\mathcal{G}\sqrt{S}$, it is possible to adiabatically eliminate all these cavity modes and obtain an effective spin-only Lindblad master equation. When $\phi_{l+1}-\phi_{l}=\pi$, one can obtain Eq. (2) with $\chi=\Gamma=4\mathcal{G}^{2}/\kappa$.

### v2

@@ V.4 Tunability and string order for arbitrary spin-$S$
Note, crucially, that for any finite value of $\Delta_{e}/\chi$ and $\Delta_{b}/\chi$, $L_{\rm corr}$ remains finite and $\mathcal{O}^{(z)}_{\rm string}(\pi)$ is always non-zero (see Fig. 8(a)). This result indicates that all the states in Eq. (33) have SPT order and belong to the same phase as the spin-1 AKLT state. This agrees with Ref. 83, which argued that the parameterized MPS in Eq. (33) with $S=1/2$ serves as a path smoothly connecting spin-$1/2$ dimer and spin-$1$ AKLT states. This type of SPT order is due to the emergent SO(3) symmetry in the large-$\Omega$ limit indicated by Eq. (32). For a finite Rabi drive $\Omega$, strictly speaking, there is no true SPT order, as the string order parameter will decay to zero at infinite distance. However, for any finite system size, one can always increase $\Omega$ to realize an approximation of SPT order, where the the string order parameter remains non-zero over a distance much larger than system size (see App. D).
As we have emphasized, a key virtue of our setup is its extreme flexibility. We now show that the above construction with $S=1/2$ directly generalizes to arbitrary spin-$S$ ensembles on each site. One can show that the correlation length $L_{\rm corr}$ remains finite for the general case, i.e., all the states in the large-$\Omega$ limit lie in the same phase. We then perform an analytical calculation of the string order parameter for arbitrary spin $S$, similar to Refs. 80, 81. As shown in App. D, it is possible to separate the dependence of system parameters $\Delta_{e},\Delta_{b},\chi$ and the dependence of the angle $\varphi$,
@@
Notice that we have $\mathcal{O}^{(z)}_{\rm string}(\pi)=0$ for integer $S$ (no SPT order), and $\mathcal{O}^{(z)}_{\rm string}(\pi)>0$ for half-integer $S$ (SPT order). This result agrees with the discussions of spin-$2S$ AKLT states in Ref. 30. We show the distinct dependence of the angle $\varphi$ for different spin $S$ in Fig. 8(b).
@@ VI Conclusion and outlook
The results presented here open several avenues for further exploration. On the metrological side, distributed quantum sensing [84] represents a particularly appealing application. While explicitly analyzed for two- and four-ensemble configurations, our approach can be scaled to larger networks, potentially enabling spatially resolved field detection and improved sensitivity scaling with system size. On the state-engineering side, it is intriguing to explore dissipative stabilization schemes for more complex many-body states. For example, our approach suggests a natural route toward realizing higher-spin AKLT states by further increasing the tunability of the setup by adding additional Hamiltonian terms that have the form of rank-$1$ spherical tensors. Another interesting direction is to increase complexity by having multiple spin ensembles now with different sizes (see Ref. 85 for a related mean-field study). In addition, the stabilized states in our scheme bear a clear resemblance to quantum spin liquids [86], both consisting of superpositions of singlets with zero total angular momentum. This analogy suggests that our framework might be extendable to the dissipative preparation of spin-liquid states.
@@ Appendix A Analytic calculation in the total angular momentum basis
This result is directly relevant for the existence of the SU(2)-symmetric unitary transformation in Eq. (22). Since $B_{J,m}$ and $C_{J,m}$ are both matrix elements of spherical tensors with $k=1$ and $q=0$, they share the same $m$ dependence due to the Wigner-Eckart theorem. This is the reason why the phases $\theta_{J}$ in Eq. (22) have no $m$ dependence.
@@
Appendix B Relaxation time scale
B.1 Lower bound for two spin ensembles
Here we derive a lower bound of the relaxation time scale to a pure steady state $|\psi_{\rm ss}^{(L)}\rangle$ based on Refs. 48, 49. We define relaxation time $t_{\rm ss}$ such that the fidelity with the steady state $F(t)=\mathrm{tr}\Big(\hat{\rho}_{\rm ss}\,\hat{\rho}(t)\Big)$ satisfies $F(t)>1-\epsilon$ for $t>t_{\rm ss}$, where $\epsilon$ is a small positive number, and $\hat{\rho}_{\rm ss}=|\psi_{\rm ss}^{(L)}\rangle\langle\psi_{\rm ss}^{(L)}|$. A lower bound of $t_{\rm ss}$ is given as follows,
@@ B.1 Lower bound for two spin ensembles
In principle the lower bound derived above can be applied to any number of spin ensembles. For two spin ensembles ($L=2$), this lower bound is relatively tight, as we shown in Fig. 3(b), the optimal $t_{\rm ss}$ obtained numerically is only a constant factor away. In this case, we can further relate the lower bound of $t_{\rm ss}$ to the steady-state spin squeezing, since $\langle\hat{S}_{1}^{z}+\hat{S}_{2}^{z}\rangle_{\rm ss}=\sum_{J=0}^{2S}|c_{J,-J}|^{2}(-J)$, and $\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}=\frac{1}{2}\sum_{J=0}^{2S}|c_{J,-J}|^{2}J$. So we have
@@
Figure 9: (a) Relaxation time scale $t_{\rm ss}$ for a spin-$1/2$ chain as a function of $\chi/\Gamma$ and system size $L$. We consider the initial state as all spins pointing down, set detunings to $\Delta_{e}=0$, $\Delta_{b}/\chi=\sqrt{2}$ (the AKLT case), set Rabi drives to $\Omega/\chi=3$. Optimal $t_{\rm ss}$ can be achieved near $\chi=\Gamma$ (unidirectional propagation) as we increase $L$. (b) At $\chi=\Gamma$, one can approximate the relaxation time scale by $\Gamma t_{\rm ss}/(2\pi)\approx(\Omega/\chi)^{2}g(L)$. Except for the regime $\Delta_{b}\ll\chi$, $g(L)$ depends weakly on $L$.
B.2 Further discussions in 1D chain
We have already shown the relaxation time scale $t_{\rm ss}$ for the AKLT state (spin-$1/2$ case) in Fig. 7(c). Here we would like to provide further numerical calculations for the relaxation time scale in a 1D chain. In Fig. 9(a), we show that the relaxation time scale $t_{\rm ss}$ under the AKLT parameters ($\Delta_{e}=0,\Delta_{b}=\sqrt{2}\chi$) as a function of the ratio $\chi/\Gamma$. As we increase the system size $L$, the optimal $t_{\rm ss}$ can be achieved near $\chi=\Gamma$ (unidirectional propagation). In this case, we expect sequential relaxation, i.e. sub-ensemble $k$ will relax in a faster rate compare to sub-ensemble $l$ if $l>k$, because the information only flows in a single direction.
However, the existence of sequential relaxation in dissipative dynamics does not mean that the relaxation time scale $t_{\rm ss}$ will simply proportional to system size $L$. As shown in Fig. 9(b), we keep the same detuning pattern (see Fig. 7(a)) as the AKLT case and vary the bulk detuning $\Delta_{b}$ (we set the edge detuning $\Delta_{e}=0$). At $\chi=\Gamma$, the relaxation time scale $t_{\rm ss}$ (an upper bound for the optimal $t_{\rm ss}$) can be approximated by
$\Gamma t_{\rm ss}/(2\pi)\approx\bigg(\frac{\Omega}{\chi}\bigg)^{2}g(L).$ (58)
In the regime $\Delta_{b}\ll\chi$, where the chiral coupling dominates, we find $g(L)$ depends linearly in $L$. While in the regime $\Delta_{b}\gg\chi$, where the chiral coupling is not playing the dominant role, we find $g(L)$ is independent of $L$. The AKLT case lies in between these two regimes and $g(L)$ depends weakly in $L$. We believe this occurs due to the collective nature of the chiral spin exchange coupling, which means that the spins can interact regardless of their distance.
It is worth to mention that the relaxation time scale is nearly the same before and after detuning permutation in a 1D chain, up to transient physics due to different overlap between the initial state and the Liouvillian eigenstates. This is because the detuning permutation is equivalent to a unitary transformation to the Lindblad master equation (see Eq. (22)) and the Liouvillian spectrum is conserved.
@@ D.1 MPS representation of 1D chain
One can also conclude that the matrix $\mathbf{A}^{Jm}$ is a spherical tensor $\hat{T}^{(k)}_{q}$ with $k=J$ and $q=m$ based on the Wigner-Eckart theorem, which is acting on the bond dimension (spin-$S$ particle), i.e., $\sum_{\alpha,\beta}[\mathbf{A}^{Jm}]_{\alpha\beta}|S,\alpha\rangle\langle S,\beta|$. The same structure is used in Ref. 72 to define the valence bond solid (VBS) state.
@@ D.2 String order parameter
Similar to the calculation in Refs. 80, 81, here we would like to evaluate the string order parameter for Eq. (80) in an infinite 1D chain,
@@ D.3 Correlation length
Here, we consider the case of $S=1/2$ and set $\varphi=\pi$. We also set $\Delta_{e}=0$ and $\Delta_{b}=\sqrt{2}\,\chi$ (parameters for spin-$1$ AKLT state in the large-Rabi-drive limit). In Fig. 10(a), a numerical calculation of $L_{\rm string}$ shows that $L_{\rm string}\propto(\Omega/\chi)^{2}$. When $L_{\rm string}\gg L$, we effectively realize the SPT phase since we cannot tell the difference in a finite-size system. One can also interpret $L/L_{\rm string}$ as a measure of the fidelity to the spin-$1$ AKLT state, as shown in Fig. 10(b). As we increase $L$, we find that $L/L_{\rm string}$ and the fidelity collapse into a single line. For example, $90\%$ fidelity corresponds to $L/L_{\rm string}\approx 0.2$. Based on the discussions in the main text, we have the relaxation time scale $t_{\rm ss}\propto(\Omega/\chi)^{2}g(L)$, leading to the AKLT relaxation time scale $t_{\rm AKLT}\propto Lg(L)$ if fixing the same fidelity to the spin-$1$ AKLT state.
@@ F.2 Many spin ensembles
Assuming the Lindbladian is diagonalizable, we can extend the uniqueness of the steady-state solution to the case of $\chi\neq\Gamma$ if the permutation symmetry between ensembles is completely broken. We also numerically confirm the uniqueness for the case of $\chi\neq\Gamma$ for small system sizes (up to $L=10$ for $S=1/2$).
@@ G.1 Raman-coupled spin-exchange interactions
One way to implement the chiral spin-exchange couplings is to engineer Raman-coupled spin-exchange interactions similar to the setup in Ref. [24]. As shown in Fig. 13(a), we define $|\uparrow\rangle$ and $|\downarrow\rangle$ states in the ground manifold with transition frequency $\omega_{0}$ (generated by either hyperfine shift or Zeeman shift from a uniform magnetic field depending on specific platforms), and excited state $|e\rangle$ with transition frequency $\omega_{e}$ with respect to $|\uparrow\rangle$ state. We then apply a magnetic-field gradient along the cavity such that the transition frequency between $|\uparrow\rangle$ and $|\downarrow\rangle$ states becomes $\omega_{0,l}=\omega_{0}+(l-1)(\Delta\omega_{0})$. To engineer chiral spin-exchange couplings between nearest-neighbor ensembles (purple arrows), we apply a two-tone drive to transition from $|\uparrow\rangle$ to $|e\rangle$ off-resonantly (solid and dashed arrows). The two tones have Rabi frequency $\Omega^{\prime}$ and $i\Omega^{\prime}$, and drive frequency $\omega_{d}$ and $\omega_{d}-\Delta\omega_{0}$ respectively. We still use the same detuning $\delta_{e}=\omega_{d}-\omega_{e}$ to the excited state $|e\rangle$ assuming $\delta_{e}$ much larger than Zeeman shifts. The cavity mode couples the transition from $|\downarrow\rangle$ to $|e\rangle$ with cavity resonant frequency $\omega_{c}$. After adiabatic elimination of the excited states ($\delta_{e}\gg\Omega^{\prime}$), the Hamiltonian is given by
@@
The second term in Eq. (121) is the chiral spin-exchange coupling between nearest-neighbor ensembles, which is due to frequency selection via the two-tone frequency difference $\Delta\omega_{0}$. Other chiral spin-exchange coupling terms can be engineered by two-tone frequency differences $2\Delta\omega_{0},3\Delta\omega_{0},\cdots$. One can also cancel the first term in Eq. (121) using a single-tone drive that generates $\sum_{l}\hat{S}^{+}_{l}\hat{S}^{-}_{l}$ with opposite $\delta_{c}$.
@@ G.2 Chiral quantum networks
Another way to implement the chiral spin-exchange couplings is to couple each spin ensemble with an optical cavity, and then connect all these cavities by a unidirectional waveguide as shown in Fig. 13(b). Based on Refs. [21, 23], the Lindblad master equation of this system can be written as
@@
$\hat{H}_{l}=\frac{\Omega}{2}(e^{i\phi^{\prime}_{l}}\hat{S}^{+}_{l}+e^{-i\phi^{\prime}_{l}}\hat{S}^{-}_{l})+\delta_{l}\hat{S}^{z}_{l}+\mathcal{G}(\hat{a}^{{\dagger}}_{l}\hat{S}^{-}_{l}+\hat{S}^{+}_{l}\hat{a}_{l}),$ (123)
where $2\mathcal{G}$ is the single-atom vacuum Rabi splitting of the cavity, $\delta_{l}$ are the drive detunings to the optical transition of the $l$-th sub-ensemble, and $\phi^{\prime}_{l}$ are the phases of Rabi drives for each sub-ensemble. We perform the following gauge transformation, $\hat{S}^{+}_{l}\rightarrow\hat{S}^{+}_{l}e^{-i\phi^{\prime}_{l}}$, $\hat{S}^{-}_{l}\rightarrow\hat{S}^{-}_{l}e^{i\phi^{\prime}_{l}}$, $\hat{a}_{l}\rightarrow\hat{a}_{l}e^{i\phi^{\prime}_{l}}$, $\hat{a}^{{\dagger}}_{l}\rightarrow\hat{a}^{{\dagger}}_{l}e^{-i\phi^{\prime}_{l}}$, Eq. (122) and (123) become
@@
$\begin{gathered}\frac{d}{dt}\hat{\rho}=-i[\hat{H}^{\prime},\hat{\rho}]+\kappa\mathcal{D}\bigg[\sum_{l}e^{-i\phi_{l}}\hat{a}_{l}\bigg]\hat{\rho},\\ \hat{H}^{\prime}=\sum_{l}\hat{H}^{\prime}_{l}+i\frac{\kappa}{2}\sum_{l>k}\Big(e^{i(\phi_{k}-\phi_{l})}\hat{a}^{{\dagger}}_{k}\hat{a}_{l}-e^{-i(\phi_{k}-\phi_{l})}\hat{a}_{k}\hat{a}^{{\dagger}}_{l}\Big),\\ \hat{H}^{\prime}_{l}=\Omega\hat{S}^{x}_{l}+\delta_{l}\hat{S}^{z}_{l}+\mathcal{G}(\hat{a}^{{\dagger}}_{l}\hat{S}^{-}_{l}+\hat{S}^{+}_{l}\hat{a}_{l}),\end{gathered}$ (124)
@@
where $\tilde{\phi}_{l}=\phi_{l}-\phi^{\prime}_{l}$. When $\kappa\gg\mathcal{G}\sqrt{S}$, it is possible to adiabatically eliminate all these cavity modes and obtain an effective spin-only Lindblad master equation. When $\tilde{\phi}_{l+1}-\tilde{\phi}_{l}=\pi$, one can obtain Eq. (2) with $\chi=\Gamma=4\mathcal{G}^{2}/\kappa$. This phase requirement can be satisfied by controlling either the separation distance between cavities holding the sub-ensembles or the phases of the Rabi drives.
Appendix H Discussions of experimental imperfections
Figure 14: (a) Steady state spin squeezing $\xi^{2}_{\rm ss}$ considering atom number imbalance between two spin ensembles. We fix the atom number in the first ensemble $N_{1}=40$, and vary the atom number in the second ensemble $N_{2}$. When $N_{1}\neq N_{2}$, $\xi^{2}_{\rm ss}$ reaches a optimal value at non-zero $\Delta/\Omega$ ratio. $\xi^{2}_{\rm ss}$ degrades as we increases the atom number imbalance. (b) Scaling of optimal steady state spin squeezing $\xi^{2}_{\rm opt}$ (optimized over $\Delta/\Omega$). We compare the cases with atom number imbalance scales as a constant value and those with imbalance scales like $\sqrt{N}$.
H.1 Population imbalance between spin ensembles
In the main text, we assume each spin ensemble has the same number of atoms. Here we analyze the effects on the steady-state spin squeezing generated by atom-number imbalance between two spin ensembles, as shown in Fig. 14. We define the atom number in the first ensemble and the second ensemble as $N_{1}$ and $N_{2}$ respectively, and define averaged atom number $N=(N_{1}+N_{2})/2$ and averaged spin value as $S=N/2$. In the presence of atom-number imbalance, the lowest total angular momentum is $J_{\rm min}=|N_{1}-N_{2}|/2>0$. As shown in Fig. 2(c), the distribution of the wave function across the different total-angular-momentum sectors depends on the ratio $\Delta/\Omega$. This effect determines the behavior of the steady state spin squeezing parameter $\xi_{\rm ss}^{2}$ (defined in Eq. (12)):
  • In the regime $S\Delta/\Omega\gg J_{\rm min}$, the steady state is close to a pure state since the relevant angular-momentum structure is very similar to balanced case $N_{1}=N_{2}$: The state populates subspaces of large total angular momentum and is not sensitive to the fact that the lowest total-angular-momentum sector has been moved from $J_{\mathrm{min}}=0$ to $J_{\mathrm{min}}=|N_{1}-N_{2}|/2$. In this regime, spin squeezing enhances as we decrease the ratio $\Delta/\Omega$, similar to Fig. 2(d).
  • In the regime $S\Delta/\Omega\ll J_{\rm min}$, the steady state mainly occupies low total-angular-momentum subspaces and becomes sensitive to the effect of population imbalance. Now, the steady state becomes a highly mixed state limited by the population imbalance, since we cannot find angular momentum smaller than $J_{\rm min}$. This causes the squeezing to degrade.
The optimal spin squeezing is typically achieved when $S\Delta/\Omega\sim J_{\rm min}$, where the steady state is roughly a pure state centered near total angular momentum $J_{\rm min}$. The qualitative understanding above agrees with our numerical findings in Fig. 14(a), in which the steady state spin squeezing $\xi^{2}_{\rm ss}$ reaches an optimal value at non-zero $\Delta/\Omega$ ratio when $N_{1}\neq N_{2}$. In Fig. 14(b), we show that the steady-state spin squeezing is mainly set by $J_{\rm min}$, i.e. $\xi^{2}_{\rm opt}\sim J_{\rm min}/N$. When $J_{\rm min}\sim\sqrt{N}$, we have $\xi^{2}_{\rm opt}\propto 1/\sqrt{N}$; When $J_{\rm min}$ is a constant value, we have $\xi^{2}_{\rm opt}\propto 1/N$.
H.2 Measurement noise
Figure 15: Classical Fisher information (CFI) $\tilde{F}_{\phi}$ for differential phase $\phi$ (near $\phi=\pi/4$) including measurement noise $\sigma_{m}$ (see text). $\sigma_{\rm SQL}=1/\sqrt{8S}$ is the SQL-level noise for excitation fractions. The dashed lines mark the spin-projection noise for excitation fractions $\sigma_{\rm QPN}$ (see text) with colors indicating the value of $\Delta/\Omega$.
Here we examine the effects of measurement noise in the case of ellipse fitting for differential phase $\phi$. The measurement noise turns the projective measurement of each spin ensemble in the basis $|S,m_{1}\rangle\otimes|S,m_{2}\rangle$ into the following POVM-type measurements,
$\hat{E}_{\tilde{p}_{1}\tilde{p}_{2}}=\bigotimes_{j=1,2}\bigg[\sum_{m_{j}}|S,m_{j}\rangle\langle S,m_{j}|\,\frac{e^{-(\tilde{p}_{j}-p_{j})^{2}/(2\sigma_{m}^{2})}}{\sqrt{2\pi\sigma_{m}^{2}}}\bigg],$ (125)
with $\int d\tilde{p}_{1}d\tilde{p}_{2}\hat{E}_{\tilde{p}_{1}\tilde{p}_{2}}=1$. Here, $p_{j}=m_{j}/(2S)+1/2$ are the excitation fractions without measurement noise, $\tilde{p}_{j}$ are the excitation fractions in the presence of measurement noise, and $\sigma_{m}$ controls the level of measurement noise. Without measurement noise ($\sigma_{m}\to 0$), we use the conditional probability $P(p_{1},p_{2}|\phi)$ defined in Eq. (18) to estimate the classical Fisher information (CFI) $F_{\phi}$ for the differential phase $\phi$ (see Eq. (19)). In the presence of measurement noise, we calculate the conditional probability using the above POVM and find
$P(\tilde{p}_{1},\tilde{p}_{2}|\phi)=\sum_{p_{1},p_{2}}P(p_{1},p_{2}|\phi)\frac{1}{2\pi\sigma_{m}^{2}}e^{-\sum_{j}(\tilde{p}_{j}-p_{j})^{2}/(2\sigma_{m}^{2})}.$ (126)
Using $P(\tilde{p}_{1},\tilde{p}_{2}|\phi)$, one can now estimate the CFI $\tilde{F}_{\phi}$ including measurement noise. In Fig. 15, we show $\tilde{F}_{\phi}$ as a function of the measurement noise $\sigma_{m}$. The full power of the quantum enhancement can only obtained in the regime $\sigma_{m}\ll\sigma_{\rm QPN}$, i.e., if the level of measurement noise is well below the spin-projection noise. Here the spin-projection noise is given by $\sigma_{\rm QPN}=\sqrt{\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}}/(4S)$, which can be interpreted as the half-thickness of the ellipse.
Note that this result is expected for all spin-squeezing protocols that do not use signal amplification before readout. If the measurement noise in an experiment is larger than the spin-projection noise, the signal-to-noise ratio is dominated by the ratio between signal and measurement noise. In this case, amplification can be used to increase the signal and the spin-projection noise until measurement noise and spin-projection noise become comparable, which improves the overall signal-to-noise ratio. For unitary OAT and dissipative single-mode spin squeezing protocols, simple signal amplification protocols have been proposed to mitigate measurement noise [89, 90]. It is an interesting topic to study how these concepts can be generalized to multimode spin squeezing and ellipse-fitting measurements.